Hello out there,I need some help with some problems,the first one is,A rectangle has side length 30 mm and diagonal length 34 mm.Find the length of the other side of the rectangle to the nearest whole number,for that problem,I;m coming up with three answers all the time,45,16,and 7.The next problem is,A right triangle has legs of lengths 10 and 24.What is the length of the hypotenuse?Round to the nearest tenths,for that one I got 26.0,is that correct?The next one is,A right triangle has a hypotenuse equal to 35 and one leg of length 30.Find the length of the other leg to the nearest tenth,and for that one I got three different answers as well,46.1,18.0,and 2.2.And the last one is,Estimate the value of the square root:0.85,and for that one I got four different answers,3,2,0,and 1,thanks for whoever can help.

Answer by stanbon(75887)

A rectangle has side length 30 mm and diagonal length 34 mm.Find the length of the other side of the rectangle to the nearest whole number,for that problem,I;m coming up with three answers all the time,45,16,and 7.
---------
Let the other side be "x".
Use Pythagoras to get:
x^2+30^2 = 34^2
x^2 = 34^2-30^2
x^2 = 256
x=16 mm
------------------
The next problem is,A right triangle has legs of lengths 10 and 24.What is the length of the hypotenuse?Round to the nearest tenths,for that one I got 26.0,is that correct?
-----------
Let the hypotenuse be "x".
x^2 = 10^2 + 24^2
x^2 = 676
x = 26.0
------------



The next one is,A right triangle has a hypotenuse equal to 35 and one leg of length 30.Find the length of the other leg to the nearest tenth,and for that one I got three different answers as well,46.1,18.0,and 2.2.
---------------
x^2+30^2 = 35^2
x^2 = 35^2 -30^2
x^2 = 325
x = 18.0
-----------------

And the last one is,Estimate the value of the square root:0.85,and for that one I got four different answers,3,2,0,and 1.
sqrt0.85 is approx 0.92
Because (0.92)^2 = 0.8464
========================
Cheers,
Stan H.